A circular pond will be placed on a square piece of land. The length of a side of the square is 2x. The radius of the pond is x. The part of the square not covered by the pond will be planted with flowers. What is the area of the region that will be planted with flowers? Write your answer in factored form.
Answer by edjones(7569)

A(sq)=s^2=(2x)^2
A(cir)=pi*r^2=pi*x^2
A(flowers)=(2x^2)-pi*x^2
=4x^2-pi*x^2
=.8584x^2
.
Ed